OpenStudy (wwe123):

find the condition that the ratio between the roots of the ax^2+bx+c=0 may me m:n

OpenStudy (saifoo.khan):

@satellite73

OpenStudy (anonymous):

i don't understand the question

OpenStudy (wwe123):

i am also

OpenStudy (wwe123):

i min

OpenStudy (anonymous):

it's simple really. there's a ratio. it's between roots. sprouting there innocently. conditionally, I may m is to n as... get to work!

OpenStudy (wwe123):

find the condition that the ratio between the roots of the equation ax^2+bx+c=0 may me m:n

OpenStudy (anonymous):

didn't quite get that @wwe123 ... could you repost it?

OpenStudy (wwe123):

can u solve this more @Algebraic!

OpenStudy (anonymous):

we know that if the roots are \(r_1\) and \(r_2\) then \(r_1r_2=\frac{c}{a}\) but somehow i don't think that answers the question \(m\) and \(n\) could be anything. are you suppossed to write \(m\) and \(n\) in terms of \(a,b,c\) or are you supposed to write \(a, b, c\) in terms of \(m,n\) ?

OpenStudy (anonymous):

I certainly couldn't solve it any less, so I guess, yes, I can solve it more.

OpenStudy (anonymous):

that is why they are called lessons

OpenStudy (anonymous):

\[\frac{m}{n}=\frac{r_1}{r_2}\] you can find \(r_1,r_2\) using the quadratic formula if nothing else comes to mind

OpenStudy (wwe123):

\[\frac{ m }{ n }=\frac{ -b+\sqrt{b ^{2}-4ac} }{-b-\sqrt{b ^{2}-4ac}}\]

OpenStudy (wwe123):

@satellite73

OpenStudy (anonymous):

or vice versa

hartnn (hartnn):

Let the roots be my & ny Therefore sum of roots ---> my+ny = -b/a \((m+n)y = -b/a\) .......(1) Product of roots---> my x ny = c/a \(mny^2=c/a \).........(2) Squaring (1), we get \((m+n)^2 y^2 = (-b/a )^2 \)...........(3) Dividing (3) by (2)…………(y^2 will be cancelled) (m+n)^2/mn = b^2/a^2 x a/c \(ac(m+n)^2 = b^2mn.\) Hence, the required condition is \(ac(m+n)^2 = b^2mn.\)

OpenStudy (wwe123):

thanks @hartnn

hartnn (hartnn):

welcome ^_^ ask if u have any doubts in any step.

OpenStudy (wwe123):

i use this method already buy who u show that Let the roots be my & ny @hartnn

hartnn (hartnn):

u mean *why? since the roots are in ratio m:n, they are in ratio of any multiples of m and n, like if the roots are in ratio 2:3 then they are in ratio 4:6 also

OpenStudy (wwe123):

ok